can you explain to me how to do the problem : Find a number whose product with 9 is the same as its usm with 56? step by step
Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

9n = n+56
8n = 56
n = 7

Answer by mananth(16946)

let number be x
product with 9 = 9x
..
sum with 56 = x+56
..
9x=x+56
9x-x=56
8x=56
x= 7
...
m.ananth@hotmail.ca